# Remote Cooling of Spin-ensembles through a Spin-mechanical Hybrid Interface

### v2

@@ Appendix A Simulation details
The coupling parameters are also $g_{0}=g=10$ kHz, with an oscillator frequency of $\omega=1.2$ MHz. Accordingly, we set the number of inversion pulses for each round of probe projection to $n=120$, see the probe control sequence in Fig. 11(b). This is to ensure that the effective probe-oscillator coupling is comparable to the oscillator frequency, as discussed in Eq. (11).
@@
Appendix B Effect of decoherence
Our numerical simulations have not considered the direct effect of spin dephasing noise. For the hybrid system considered in this work, such noise is induced by the spin bath surrounding the probe and the ensemble, and is expected to be the dominant noise source for the hybrid system considered [52]. In each measurement, both the ensemble spins and the spin bath surrounding them collapse partially and simultaneously towards the steady states. This collapse is sensitive only to the probe population in the measurement basis [53].
Similarly, in the case of the oscillator, its dissipation ($T_{1}$) also has a negligible effect. As shown in Fig. 4, the cooling of the spin ensemble requires more measurements and time compared to the cooling of the oscillator. Since the (initial) state of the oscillator is unaffected during the cooling of the ensemble, the effect of the oscillator’s $T_{1}$ can be ignored. Counterintuitively, in some cases rapid decoherence of the intra-bath correlations can actually increase the cooling rate rather than decrease it [54]. See further discussion of the effect of oscillator $T_{1}$ decoherence (relaxation) on its cooling in our previous work [43].
Appendix C Encoding a qubit into a macroscopic oscillator
@@ B Experimental feasibility
From an experimental point of view, it’s crucial that the coherence time of the oscillator is considerably long compared to the coupling strength $g_{0}$, ideally at least on the order of $100/g_{0}$. Diamond NMOs with a quality factor $Q>10^{6}$ have been demonstrated at sub-Kelvin temperatures and mechanical damping rates of a few kHz [59]. This implies that the coherence time of the vibrational mode must be significantly improved. Note that the coherence time of the NV electronic spin in cryogenic conditions can exceed 1 second by the application of dynamical decoupling [52]. This is far from sufficient, as the probe spin only needs to maintain coherence during each phase estimation round.

### v3

@@ Appendix C Simulation details
For the simulations conducted in the main text, the coupling parameters are also $g_{0}=g=10$ kHz, with an oscillator frequency of $\omega=1.2$ MHz. Accordingly, we set the number of inversion pulses for each round of probe projection to $n=120$, see the probe control sequence in Fig. 11(b). This is to ensure that the effective probe-oscillator coupling is comparable to the oscillator frequency, as discussed in Eq. (11).
@@
Appendix D Encoding a qubit into a macroscopic oscillator
@@ B Experimental feasibility
From an experimental point of view, it’s crucial that the coherence time of the oscillator is considerably long compared to the coupling strength $g_{0}$, ideally at least on the order of $100/g_{0}$. Diamond NMOs with a quality factor $Q>10^{6}$ have been demonstrated at sub-Kelvin temperatures and mechanical damping rates of a few kHz [58]. This implies that the coherence time of the vibrational mode must be significantly improved. Note that the coherence time of the NV electronic spin in cryogenic conditions can exceed 1 second by the application of dynamical decoupling [39]. This is more than sufficient, as the probe spin only needs to maintain coherence during each phase estimation round.